# Disentangling Brillouin’s negentropy law of information and Landauer’s law on data erasure

### v2

@@ Abstract
The link between information and energy introduces the observer and their knowledge into the understanding of a fundamental quantity of physics. Two approaches compete to account for this link, Brillouin’s negentropy law of information and Landauer’s law on data erasure, which are often confused. The first, based on the Clausius’ inequality and Shannon’s mathematical results is very robust, while the second, based on the simple idea that information needs a material embodiment (data-bits) is today perceived as more physical and prevails. In this paper, we show that Landauer’s idea results from a confusion between information (a global emergent concept) and data (a local material object). This confusion leads to many inconsistencies and is incompatible with thermodynamics and information theory. The reason it prevails is interpreted to be due to a frequent tendency of materialism towards reductionism, neglecting emergence and seeking to eliminate the role of the observer. A paradoxical trend given that it is often accompanied by the materialist idea that all scientific knowledge nevertheless originates from observation. Information and entropy are actually emergent quantities introduced in the theory by convention.
@@ I Introduction
The definition of something is a statement that allows us to recognize it when we see it. For example, it can be the statement of all the characteristics of that thing. Or, it can be the statement that this thing is the name given to a category of elements already defined. Of course, this supposes that the list of all elements in this category is known, otherwise we would be unable to recognize every time an element as belonging to this category.
@@
As a working definition, one can say that “in the everyday world outside of the formal language of physics, energy is the ability to do anything… Insofar as “does” means “produces a change in what was before”, energy is philosophically the determinant of all observable change”(E. Hecht) [2]). With this qualitative definition, there is no doubt that knowledge and information would be a priori forms of energy. An example is given by Maxwell [3]: an intelligent being (a demon) produces mechanical work from the information they get about a device. If energy cannot come from nowhere according to the principle of conservation, we see that, a posteriori as well, information is energy.
@@
How to be quantitative about information and its conversion into unit of energy? The solution was given by Shannon’s information theory [4] which, by the intermediate of Gibbs statistical mechanics [5], allowed us to identify the entropy of a system as the uncertainty we have about it. Hence the Brillouin’s law of information [6, 7]: decreasing the uncertainty by acquiring information has a minimum energy cost equal to that of the corresponding decrease in entropy.
@@
Is that enough? No, according to Landauer. Energy is a physical quantity, information must be too [8, 9, 10]. But Landauer means by physical, something that is not abstract but tangible and material, in other word an hardware set of data-bits. Being materialized, information (understand data-bits) must be erased prior to acquisition. Hence, the Landauer’s idea: the energy cost for the acquisition of information is paid by this preliminary erasure step. This is the Landauer’s law on data-erasure.
@@
Here, we will see that what may appear as an addition or improvement to the triplet (thermodynamics-statistical mechanics-information theory) is in reality totally incompatible with it. The abstract concept of information cannot be confused with its material support (data-bit), without introducing many inconsistencies into this theory.
@@
The article is organized as follows. The first part provides a brief overview of this threefold theory (thermostatistics) which in fact forms a whole. Particular attention will be paid to the role assigned to the observer and to conventions. Next, the link between information and energy as established by the demonic engines is presented, with Brillouin and Landauer’s views on how they work. Both are often confused (e.g.[11, 12]), so that this part aims in particular to clarify their differences.
@@
Energy is a concept related to the physical world, while information refers to us as intelligent beings endowed with a mind. Even materialists (like myself) have to force themselves to connect the two. That is why, in my opinion, it is an illusion to think that the different conceptions of what science is, of what a theory is, play no role in this problem. Even though physicists often wish to distance themselves from philosophy, an epistemological approach is necessary to understand the interplay between information, energy, probabilities and the role of the observer. Epistemological positions are always present in the literature on this subject, and in particular in that of Landauer. In most cases, they are hidden or implicit. This does not favor clarity.
@@
We also adopt an epistemological position throughout this article, but it will be explicit from the outset. It will be that of neo-positivists [13, 14, 15, 16] and those who influenced them [17, 18, 19, 20]. In brief: 1) only two sources of knowledge are admitted, logic and observation; 2) an explanation is a deduction of observations (past and future) from theory; 3) the theory includes inductive laws (generalization of observations), but it must necessarily also include conventions chosen according to their convenience and to the economy of thought they allow. The last part of the paper shows how this position permits to analyze Landauer’s reductionist idea (beyond its inconsistencies) and how to include emergentism into materialism.
@@ II.1 Thermodynamics: principle versus law
Energy is one of the few concepts in physics with universal scope. So that every scientist knows the fundamental statements on which the theory of thermodynamics, the science of energy transformations, is based. In short:
@@
1st principle: The total quantity of energy of an isolated system (the universe) is conserved during any change.
@@
2nd law: There exists a state quantity, which we call entropy $S$, such that for any change from state A to state B, $\Delta S_{\text{AB}}=(S_{\text{B}}-S_{\text{A}})\geq\int\mathop{}\!\mathrm{d}Q/T$ (Clausius inequality), where $Q$ and $T$ are the heat exchanged and temperature, respectively.
@@
But what people are less aware of is the fundamental difference in nature between the last two statements, hence the terms “principle” and “law” used to highlight this point (this denomination follows that of Poincaré [18]). A principle is a convention, a law is inferred from induction. Let us clarify this point starting with the second law.
@@
Clausius inequality expresses that, for a given change of state along a differentiable path, the heat (in temperature unit) received by the system from the environment never exceeds a certain limit that depends solely on the initial and final states. This limit is the difference in entropy. It is usually approached as the rate of change along the trajectory is slowed down in a quasi-static manner. The paths that allow this are called reversible because they can be followed in both directions. The others, especially those which are not differentiable, are called irreversible. In Clausius inequality, the sign of $Q$ refers to the system (positive when received). For simplicity, consider a system with constant internal energy, typically a set of independent entities, such as an ideal gas or a set of bits at a given temperature. Then for any change, transfers of mechanical work ($W$) and heat ($Q$) with the environment compensate each other ($\mathop{}\!\mathrm{d}Q=-\mathop{}\!\mathrm{d}W$), so that the Clausius inequality rewrites by considering the reverse change from B to A: $\Delta S_{\text{BA}}\leq\int\mathop{}\!\mathrm{d}W/T$. It follows that the entropy difference essentially corresponds to the observed minimum work (in temperature unit) required for the system to return to its initial state by a quasi-static path (hence the word “entropy” from Greek entropia “a turning toward”). It is clear that mechanical work and heat can be measured in as many different circumstances as one can imagine. The Clausius inequality results from the generalization of these specific observations, the regularity of which has never been contradicted in two centuries of experiments in this area. The case of equality allows us to define the word “entropy”, but this is just a nominal definition, the core of the second law lies in the inequality. The second law follows from induction.
@@
As for the first principle, it states that the energy of the universe is conserved regardless of the change undergone by a system. In other words, the energy balance is always zero. One could imagine that the first principle also derives from precise measurements of the energy balance of various changes, which would lead to the conclusion that it is apparently always zero, thus allowing these observations to be generalized and elevated to the rank of a universal law. Of course, this assumes that we are able to recognize (identify) energy, whatever form it takes, when we see it. But that is not how it happened, because there is absolutely nothing that allows us to recognize something a priori as being a form of energy before this form has been added to the list of those known. And this is precisely this kind of measurements that allows for such an addition. The first principle does not follow from inductive reasoning. It is actually the hidden, but very incomplete, definition of what energy is. Incomplete because if a quantity is conserved, meaning constant over time, any function of this quantity is also conserved and constant over time. So that, if a quantity is conserved, many others are also conserved. Among them, what should we understand by ”energy”? Despite its incomplete nature, we must be content with this definition because there is no other [1].
@@
Although this would be extremely unlikely for the second law, there is no conceptual impediment to discovering a counterexample that would invalidate the generality of any inductive law. But for a definition, such as the first law, such an invalidation is conceptually meaningless. A definition cannot be invalidated by an observation: “[A] principle…is no longer subject to the test of experiment. It is not true or false, it is convenient”[19]. In case an experiment would lead to a non-zero energy balance, we have two options: 1) either we decide that the 1st principle is no longer appropriate; 2) or we acknowledge that something escaped to us and that we have just discovered a new form of energy. Clearly, the second option is more convenient, and that is how all forms of energy were added to the list of those known [14], for instance kinetic energy (energy of motion) or rest mass (energy at rest). Ultimately, energy is defined only as a category comprising an undetermined number of items, a number that depends on the state of the art. Energy is an abstract conventional concept, it is like a universal in the metaphysical sense given to this word. But this universal has a particularity: it includes a number of elements that depends on our knowledge. So that internal energy cannot be considered as intrinsic to the state of a system, nor as intrinsic to certain phenomena. That is to say, energy cannot be considered independent of the observer. “Energy is […] the determinant of all observable change” (E. Hecht [2]).
@@
What about entropy? The notion of change is central in the first principle defining energy (energy is conserved when everything else change), but also in the second law defining entropy. The latter is not defined in an absolute way by the Clausius’ inequality, but only in a relative way from the measure of the heat exchanged when the system undergoes a change of state. No observable change, no difference of entropy. We could say the same for any state quantity, for example temperature. But for temperature, we have means to measure its value for a given state without the need of any change. For entropy there is no such mean. Entropy is a state quantity that can only be deduced from the observation of a change of state. Ultimately, entropy depends on the ability of the observer to observe these changes, that is to say from the ability of the observer to perceive differences. No work is needed to return to the initial state if no change is observed: “The idea of dissipation of energy depends on the extent of our knowledge” (J. C. Maxwell [21]).
@@ II.2 Statistical mechanics: reductionism versus emergentism
Statistical mechanics, “the rational foundation of thermodynamics”[5], aims to remedy the subjectivity introduced by the observer in thermodynamics. The goal is to calculate everything from the Newton’s mechanics of atoms, that is to say from the behavior of objects totally independent of us. It is a reductionist approach.
At the atomic scale everything is in motion. Microscopic configurations (also named phases or microstates) are constantly changing. But they are also so numerous that it is impossible to hope to know that of a system other than in terms of probabilities. So, statistical mechanics faces two main problems:
@@
    Prior distribution: What phase probability distribution should be used as a starting point for calculations, given that it cannot be measured?
@@
The first problem is generally solved by the ergodic hypothesis, which essentially considers that the phases distribution of snapshots of identical systems (which differ only in their phase) is equal to that of a single system evolving in time. The trajectory of the system in the phase space $\Gamma$ is supposed to be volume preserving, like that of a deterministic mechanical system (Liouville’s theorem). For two successive points $a$ and $b$ of the trajectory, if the system is deterministic the probability for the system to be in $b$ knowing it was in $a$ is equal to 1. So that the probability for $b$ is equal to that of $a$, and this is true throughout the trajectory. This results in a uniform probability distribution of phases for isolated systems, thus opening the door to additional calculations and great results. In particular the Boltzmann distribution for a closed system and the temperature dependence of the partition function [22], which by identification to known equations of thermodynamics leads finally to the famous Gibbs’ result for entropy:
@@
Clausius entropy, initially defined in thermodynamics solely by its measurable variations, was equal to the minimum amount of heat received by the environment. Thanks to statistical mechanics, it is now understood in absolute terms as the logarithm of the number of possibilities for the microscopic configuration. If understanding means creating connections, then progress is huge.
But getting over the observer is not without introducing new problems. Where has the observer’s knowledge gone, the information he possesses? The best illustration of this problem is given by the famous two Gibbs paradoxes [23, 24]:
@@
    Mixing two volumes of gas requires work to return to the initial state only if these two gases were initially identified as different. But for statistical mechanics, both cases increases entropy because replacing the partition between the two volumes is not enough to ensure that each particle returns to its original compartment.
@@
Actually in these two problems, statistical mechanics consider the overall information needed to describe the system (which particle in which compartment), whereas thermodynamics consider an incomplete information. “It is to states of systems thus incompletely defined that the problems of thermodynamics relate” (J.W. Gibbs  [25]).
@@
In thermodynamics, the perception of a change depends not only on the observer’s knowledge, but also on what they consider relevant. In a certain sense, it is true that “no man ever steps in the same river twice” (Heraclitus). But on the other hand, what makes the identity of the river and ours is not that of the molecules that constitute us: tomorrow, it will still be the Seine that flows through Paris and I hope to still be myself. The identity of molecules is information that exists, at least for large traceable molecules, but it is considered in thermodynamics by the observer as meaningless or irrelevant in certain cases (e.g. for open systems). The level of information considered relevant is a convention. And it is only in this way, by reintroducing the observer and the information they consider relevant, that Gibbs paradoxes can be resolved for large traceable molecules (see [26] p.13-14 and [27]).
@@
The second problem of statistical mechanics is related to the definition of equilibrium. The current phase is constantly changing. We could simply define the equilibrium as the macroscopic state whose properties take the average values of those of the microscopic configurations. But this is not enough, something is missing. From experience, I know that the equilibrium state of a gas in a room is that in which it uniformly occupies the entire volume and not just a corner. Fluctuations occur which take the system away from equilibrium, but there is something that restores it. Consider a gas inside a cylinder below a piston. We can push or pull the piston and experience a force (Figure 1). The system behaves as a metal spring would. The difference is that for the metal spring, the macroscopic equilibrium reflects the microscopic equilibrium of the atoms that are in a potential well. Such a potential well does not exist for the atoms of an ideal gas. One could argue that the equilibrium position of the piston corresponds to the equality of internal and external pressures, and calculate the pressure as being related to the number of atomic collisions on both side. In doing so, as for the metal spring, macroscopic forces would ultimately be explained by microscopic ones. However, the calculation inevitably starts with the assumption that the atom density near the surface is equal to that of the gas at equilibrium for a given piston position. In other words, the calculation starts with the very assumption it intends to prove. This is circular reasoning.
In reality, the force exerted on the piston is an emergent property, just like pressure and entropy. It cannot be derived consistently from the microscopic scale and the laws we already have at our disposal. An additional ingredient is necessary that must be postulated upstream in the theory by the conventional definition we give of macroscopic equilibrium. Just as a metal spring minimizes its potential energy at equilibrium, it would be convenient if a volume of gas would optimize something. In thermodynamics we already have the second law that says that $\mathop{}\!\mathrm{d}S\geq 0$ for an isolated system. If we postulate that entropy is maximum at equilibrium the trick is done [28] and we have the restoring force we are looking for. The equilibrium would become an attractor.
@@
This is where the problem arises with such a postulate within the framework of statistical mechanics as we have presented it so far. According to Poincaré’s recurrence theorem [29] a volume preserving dynamical system is recurrent and has no attractor. This is inherent to the ergodic hypothesis and its conception of probabilities as frequencies of occurrence: frequency implies recurrence. The solution would consist in getting out the ergodic hypothesis. But then comes back the problem of the prior probability distribution of phases. It is a vicious circle. Finally, the best approach would be to admit that “When one does not know anything the answer is simple. One is satisfied with enumerating the possible events and assigning equal probabilities to them.” (R. Balian [30]). This is known as the fundamental postulate of statistical mechanics but in reality nothing more than the Laplace’s “principle of insufficient reason”[31]. The problem is that, as it stands, it is unfounded. It is a synthetic a priori knowledge: “It cannot be that because we are ignorant of the matter we know something about it” (R.L. Ellis [32]).
Information theory provides us precisely the missing pieces that we lack. It reintroduces the observer’s knowledge (because information is meaningless without an observer) and solves the problem of prior probabilities by making the principle of insufficient reason analytical.
II.3 Information theory: come back of the observer
Shannon [4] tackled a problem that appeared very different from that of thermostatistics, but which ultimately turned out not to be so far: the lossless compression of a message. His started by getting out the message meaning and by considering its emitter as a source of a random variable taking its values among a set $\Gamma$ of possible characters. In doing so, he mathematically demonstrated two results:

### v3

@@ Abstract
The link between information and energy introduces the observer and their knowledge into the understanding of a fundamental quantity in physics. Two approaches compete to account for this link—Brillouin’s negentropy law of information and Landauer’s law on data erasure—which are often confused. The first, based on Clausius’ inequality and Shannon’s mathematical results, is very robust, whereas the second, based on the simple idea that information requires a material embodiment (data bits), is now perceived as more physical and therefore prevails. In this paper, we show that Landauer’s idea results from a confusion between information (a global emergent concept) and data (a local material object). This confusion leads to many inconsistencies and is incompatible with thermodynamics and information theory. The reason it prevails is interpreted as being due to a frequent tendency of materialism towards reductionism, neglecting emergence and seeking to eliminate the role of the observer. A paradoxical trend, considering that it is often accompanied by the materialist idea that all scientific knowledge, nevertheless, originates from observation. Information and entropy are actually emergent quantities introduced in the theory by convention.
@@ I Introduction
The definition of something is a statement that allows us to recognize it when we see it. For example, it can be a statement of all the characteristics of that thing. Or, it can be the statement that this thing is the name given to a category of elements already defined. Of course, this supposes that the list of all elements in this category is known; otherwise, we would be unable to reliably recognize something as belonging to this category.
@@
As a working definition, one can say that “in the everyday world outside of the formal language of physics, energy is the ability to do anything… Insofar as “does” means “produces a change in what was before”, energy is philosophically the determinant of all observable change” (E. Hecht) [22]). With this qualitative definition, there is no doubt that knowledge and information are a priori forms of energy. An example is provided by Maxwell [44]: an intelligent being (a demon) produces mechanical work from the information they obtain about a device. If energy cannot come from nowhere, according to the principle of conservation, we see that, a posteriori as well, information is energy.
@@
How to be quantitative about information and its conversion into units of energy? The solution was provided by Shannon’s information theory [50], which, through the intermediary of Gibbs statistical mechanics [20], allowed us to identify the entropy of a system as the uncertainty we have about it. Hence, Brillouin’s law of information [5, 6]: decreasing uncertainty by acquiring information has a minimum energy cost equal to that of the corresponding decrease in entropy.
@@
Is that enough? No, according to Landauer. Energy is a physical quantity; information must be too [37, 38, 39]. However Landauer means by physical something that is not abstract but tangible and material—in other words, a hardware set of data bits. Being materialized, information (understood as data bits) must be erased prior to acquisition. Hence, Landauer’s idea: the energy cost of acquiring information is paid through this preliminary erasure step. This is Landauer’s law on data erasure.
@@
Here, we will see that what may appear as an addition or improvement to the triplet (thermodynamics–statistical mechanics–information theory) is, in reality, totally incompatible with it. The abstract concept of information cannot be confused with its material support (data bit) without introducing many inconsistencies into this theory.
@@
The article is organized as follows. The first part provides a brief overview of this threefold theory (thermostatistics), which in fact forms a whole. Particular attention will be paid to the role assigned to the observer and to conventions. Next, the link between information and energy as established by the demonic engines is presented, together with Brillouin’s and Landauer’s views on how they work. Both are often confused (e.g., [24, 25]); thus, this part aims, in particular, to clarify their differences.
@@
Energy is a concept related to the physical world, while information refers to us as intelligent beings endowed with a mind. Even materialists (like myself) have to force themselves to connect the two. That is why, in my opinion, it is an illusion to think that the different conceptions of what science is, of what a theory is, play no role in this problem. Even though physicists often wish to distance themselves from philosophy, an epistemological approach is necessary to understand the interplay between information, energy, probabilities, and the role of the observer. Epistemological positions are always present in the literature on this subject, and in particular in that of Landauer. In most cases, they are hidden or implicit. This does not favor clarity.
@@
We also adopt an epistemological position throughout this article; however, it will be explicit from the outset. It will be that of neo-positivists [21, 41, 42, 8] and those who influenced them [12, 48, 49, 14]. In brief, (1) only two sources of knowledge are admitted: logic and observation; (2) an explanation is a deduction of observations (past and future) from theory; (3) the theory includes inductive laws (generalization of observations), but it must necessarily also include conventions chosen according to their convenience and to the economy of thought they allow. The last part of the paper shows how this position allows us to analyze Landauer’s reductionist idea (beyond its inconsistencies) and how to include emergentism into materialism.
@@ II.1 Thermodynamics: Principle Versus Law
Energy is one of the few concepts in physics with universal scope. As such, every scientist knows the fundamental statements on which the theory of thermodynamics, the science of energy transformations, is based. In short:
@@
First principle: The total quantity of energy of an isolated system (the universe) is conserved during any change.
@@
Second law: There exists a state quantity, which we call entropy $S$, such that for any change from state A to state B, $\Delta S_{\text{AB}}=(S_{\text{B}}-S_{\text{A}})\geq\int\mathop{}\!\mathrm{d}Q/T$ (Clausius inequality), where $Q$ and $T$ are the heat exchanged and the temperature, respectively.
@@
But what people are less aware of is the fundamental difference in nature between the last two statements; hence, the terms “principle” and “law” are used to highlight this point (this denomination follows that of Poincaré [48]). A principle is a convention, a law is inferred from induction. Let us clarify this point, starting with the second law.
@@
Clausius inequality expresses that, for a given change of state along a differentiable path, the heat (in temperature units) received by the system from the environment never exceeds a certain limit that depends solely on the initial and final states. This limit is the difference in entropy. It is usually approached as the rate of change along the trajectory slows down in a quasi-static manner. The paths that allow this are called reversible because they can be followed in both directions. The others, especially those that are not differentiable, are called irreversible. In the Clausius inequality, the sign of $Q$ refers to the system (positive when received). For simplicity, consider a system with constant internal energy, typically a set of independent entities, such as an ideal gas or a set of bits at a given temperature. Then, for any change, transfers of mechanical work ($W$) and heat ($Q$) with the environment compensate each other ($\mathop{}\!\mathrm{d}Q=-\mathop{}\!\mathrm{d}W$), so that the Clausius inequality rewrites by considering the reverse change from B to A: $\Delta S_{\text{BA}}\leq\int\mathop{}\!\mathrm{d}W/T$. It follows that the entropy difference essentially corresponds to the observed minimum work (in temperature units) required for the system to return to its initial state by a quasi-static path (hence the word “entropy” from Greek entropia “a turning toward”). It is clear that mechanical work and heat can be measured in as many different circumstances as one can imagine. The Clausius inequality results from the generalization of these specific observations, the regularity of which has never been contradicted in two centuries of experiments in this area. The case of equality allows us to define the word “entropy”, but this is just a nominal definition; the core of the second law lies in the inequality. The second law follows from induction.
@@
As for the first principle, it states that the energy of the universe is conserved regardless of the change undergone by a system. In other words, the energy balance is always zero. One could imagine that the first principle also derives from precise measurements of the energy balance of various changes, which would lead to the conclusion that it is apparently always zero, thus allowing these observations to be generalized and elevated to the rank of a universal law. Of course, this assumes that we are able to recognize (identify) energy, whatever form it takes, when we encounter it. However, this is not how it happened, because there is absolutely nothing that allows us to recognize something a priori as a form of energy before it has been added to the list of known forms. And this is precisely this kind of measurement of energy balance that allows for such an addition. The first principle does not follow from inductive reasoning. Rather, it is a hidden—but very incomplete—definition of what energy is. It is incomplete because if a quantity is conserved, that is, constant over time, any function of this quantity is also conserved and therefore constant over time. Therefore, if a quantity is conserved, many others are also conserved. Among them, what should we understand by “energy”? Despite its incomplete nature, we must be content with this definition because there is no other [17].
@@
Although this would be extremely unlikely for the second law, there is no conceptual impediment to discovering a counterexample that would invalidate the generality of any inductive law. However, for a definition such as the first principle, such an invalidation is conceptually meaningless. A definition cannot be invalidated by an observation: “[A] principle… is no longer subject to the test of experiment. It is not true or false, it is convenient” [49]. In case an experiment would lead to a non-zero energy balance, we would have the following two options: (1) either we decide that the first principle is no longer appropriate; or (2) we acknowledge that something has escaped our notice and that we have just discovered a new form of energy. Clearly, the second option is more convenient, and that is how all forms of energy were added to the list of those known [41]; for instance, kinetic energy (energy of motion) and rest mass (energy at rest). Ultimately, energy is defined only as a category comprising an undetermined number of items—a number that depends on the state of the art. Energy is an abstract, conventional concept; it is like a universal in the metaphysical sense attributed to this word. However, this universal has a particularity: it includes a number of elements that depend on our knowledge. Therefore, internal energy cannot be considered as intrinsic to the state of a system, nor as intrinsic to certain phenomena. That is to say, energy cannot be considered independent of the observer. “Energy is […] the determinant of all observable change” (E. Hecht [22]).
@@
What about entropy? The notion of change is central to the first principle, which defines energy—energy is conserved when everything else changes—but also to the second law, which defines entropy. The latter is not defined in an absolute way by Clausius’ inequality, but only in a relative way, through the measurement of the heat exchanged when the system undergoes a change of state. No observable change means no difference in entropy. We could say the same for any state quantity, for example, temperature. But for temperature, we have means to measure its value for a given state without the need for any change. For entropy, there is no such mean. Entropy is a state quantity that can only be deduced from the observation of a change of state. Ultimately, entropy depends on the ability of the observer to observe these changes, that is, on the observer’s ability to perceive differences. No work is needed to return to the initial state if no change is observed: “The idea of dissipation of energy depends on the extent of our knowledge” (J.C. Maxwell [45]).
@@ II.2 Statistical Mechanics: Reductionism Versus Emergentism
Statistical mechanics, “the rational foundation of thermodynamics” [20], aims to remedy the subjectivity introduced by the observer in thermodynamics. The goal is to calculate everything from Newton’s mechanics of atoms, that is, from the behavior of objects totally independent of us. It is a reductionist approach.
At the atomic scale, everything is in motion. Microscopic configurations (also named phases or microstates) are constantly changing. However, they are so numerous that it is impossible to hope to know that of a system other than in terms of probabilities. Thus, statistical mechanics faces the following two main problems:
@@
    Prior distribution: What phase probability distribution should be used as a starting point for calculations, considering that it cannot be measured?
@@
The first problem is generally solved by the ergodic hypothesis, which essentially considers that the phase distribution of snapshots of identical systems (which differ only in their phase) is equal to that of a single system evolving in time. The trajectory of the system in the phase space $\Gamma$ is supposed to be volume preserving, like that of a deterministic mechanical system (Liouville’s theorem). For two successive points $a$ and $b$ of the trajectory, if the system is deterministic, the probability for the system to be in $b$ knowing it was in $a$ is equal to 1. Thus, the probability of $b$ is equal to that of $a$, and this holds throughout the trajectory. This results in a uniform probability distribution of phases for isolated systems, thus opening the door to additional calculations and great results. In particular, the Boltzmann distribution for a closed system and the temperature dependence of the partition function [31], which—by identification with known thermodynamics equations—ultimately lead to the famous Gibbs’ result for entropy:
@@
Clausius entropy, initially defined in thermodynamics solely by its measurable variations, was equal to the minimum amount of heat received by the environment. Thanks to statistical mechanics, it is now understood in absolute terms as the logarithm of the number of possibilities for the microscopic configuration. If understanding means creating connections, then progress is substantial.
However, removing the observer is not without introducing new problems. Where has the observer’s knowledge gone, the information he possesses? The best illustration of this problem is provided by the following two Gibbs paradoxes [46, 10]:
@@
    Mixing two volumes of gas requires work to return to the initial state only if these two gases were initially identified as different. However for statistical mechanics, both cases increase entropy because replacing the partition between the two volumes is not enough to ensure that each particle returns to its original compartment.
@@
Actually, in these two problems, statistical mechanics considers the overall information needed to describe the system (which particle in which compartment), whereas thermodynamics considers incomplete information. “It is to states of systems thus incompletely defined that the problems of thermodynamics relate” (J.W. Gibbs [19]).
@@
In thermodynamics, the perception of a change depends not only on the observer’s knowledge, but also on what they consider relevant. In a certain sense, it is true that “no man ever steps in the same river twice” (Heraclitus). On the other hand, what makes the identity of the river and ours is not that of the molecules that constitute us: tomorrow, it will still be the Seine that flows through Paris, and I hope to still be myself. The identity of molecules is information that exists, at least for large traceable molecules, but it is considered in thermodynamics by the observer as meaningless or irrelevant in certain cases (e.g., for open systems). The level of information considered relevant is a convention. And it is only in this way, by reintroducing the observer and the information they consider relevant, that Gibbs paradoxes can be resolved for large traceable molecules (see [36] (pp. 13–14) and [32]).
@@
The second problem of statistical mechanics is related to the definition of equilibrium. The current phase is constantly changing. We could simply define the equilibrium as the macroscopic state whose properties take the average values of those of the microscopic configurations. However, this is not sufficient; something is missing. From experience, I know that the equilibrium state of a gas in a room is that in which it uniformly occupies the entire volume and not just a corner. Fluctuations occur that take the system away from equilibrium, but there is something that restores it. Consider a gas inside a cylinder below a piston. We can push or pull the piston and experience a force (Figure 1). The system behaves as a metal spring would. The difference is that for the metal spring, the macroscopic equilibrium reflects the microscopic equilibrium of the atoms that are in a potential well. Such a potential well does not exist for the atoms of an ideal gas. One could argue that the equilibrium position of the piston corresponds to the equality of internal and external pressures, and calculate the pressure as being related to the number of atomic collisions on both sides. In this way, as for the metal spring, macroscopic forces would ultimately be explained by microscopic ones. However, the calculation inevitably starts with the assumption that the atom density near the surface is equal to that of the gas at equilibrium for a given piston position. In other words, the calculation starts with the very assumption it intends to prove. This is circular reasoning.
@@
In reality, the force exerted on the piston is an emergent property, just like pressure and entropy. It cannot be derived consistently from the microscopic scale and the laws we already have at our disposal. An additional ingredient is necessary, which must be postulated upstream in the theory through the conventional definition we give of macroscopic equilibrium. Just as a metal spring minimizes its potential energy at equilibrium, it would be convenient if a volume of gas optimized something. In thermodynamics, we already have the second law that says that $\mathop{}\!\mathrm{d}S\geq 0$ for an isolated system. If we postulate that entropy is maximum at equilibrium, the problem is resolved [7], and the restoring force we are seeking naturally emerges. The equilibrium would become an attractor.
This is where the problem arises with such a postulate within the framework of statistical mechanics as we have presented it so far. According to Poincaré’s recurrence theorem [47], a volume-preserving dynamical system is recurrent and admits no attractor. This property is inherent to the ergodic hypothesis and to its conception of probabilities as frequencies of occurrence: frequency implies recurrence. The solution would consist in abandoning the ergodic hypothesis. However, then comes back the problem of the prior probability distribution of phases. It is a vicious circle. Finally, the best approach would be to admit that “When one does not know anything, the answer is simple. One is satisfied with enumerating the possible events and assigning equal probabilities to them.” (R. Balian [2]). This is known as the fundamental postulate of statistical mechanics, but in reality, nothing more than Laplace’s “principle of insufficient reason” [11]. The problem is that, as it stands, it is unfounded. It is a synthetic a priori knowledge: “It cannot be that because we are ignorant of the matter we know something about it” (R.L. Ellis [16]).
Information theory provides us precisely with the missing pieces that we lack. It reintroduces the observer’s knowledge (because information is meaningless without an observer) and solves the problem of prior probabilities by making the principle of insufficient reason analytical.
II.3 Information Theory: The Return of the Observer
Shannon [50] tackled a problem that initially appeared very different from that of thermostatistics, but which ultimately turned out not to be so distant: the lossless compression of a message. Shannon began by abstracting away the meaning of the message and by modeling its emitter as a source of a random variable taking values in a set $\Gamma$ of possible characters. In this framework, he mathematically demonstrated the following two results: